### v1

# Flow Graph-Based Classification of Defects4J Faults

Abstract
Software fault datasets such as Defects4J provide for each individual fault its location and repair, but do not characterize the faults. Current classifications use the repairs as proxies, which does not capture the intrinsic nature of the fault. In this paper, we propose a new, direct fault classification scheme based on the control- and data-flow graph representations of the program. Our scheme comprises six control-flow and two data-flow fault classes. We apply this to 488 faults from seven projects in the Defects4J dataset. We find that one of the data-flow fault classes (definition fault) is the most common individual class but that the majority of faults are classified with at least one control-flow fault class. The majority of the faults are assigned between one and three classes. Our proposed classification can be applied to other fault datasets and can be used to improve fault localization and automated program repair techniques for specific fault classes.
I Introduction
Software debugging is an expensive and time-consuming process [1, 2], requiring faults to be found (fault localization) and fixed (fault repair). Its effectiveness depends on several factors, but one factor that is often overlooked is the nature of the fault itself [3]. Typically, this fault nature is unknown, but an analysis of known real-world faults allows us to identify common fault types, which can in turn be used to improve fault localization and repair techniques.
However, investigating the intrinsic nature of faults can be labor-intensive, and classifications typically use the applied repairs as proxies [4, 3, 5, 2]. This allows automating the classification process but it does not necessarily capture the intrinsic nature of the faults themselves, since faults can typically be repaired in different ways.
In this paper, we describe a new fault classification scheme based on control- and data-flow graphs. The main novelty of our scheme is its grounding in the (abstract) structural properties of the faults themselves, rather than the applied repairs. This not only eliminates any potential bias by repair-based classifications, it also emphasizes the commonalities of different faults under syntactic variations. Consider for example the faults (and repairs) shown in LABEL:listing:lang62 and LABEL:listing:chart18. In both cases, the underlying problem is that the control flow from the fault location to the next location is wrong (i.e., whenever the program executes the faulty statement, it executes the wrong statement next), despite the differences in the faulty statements (i.e., return vs. missing statement) and the repair (i.e., substitution of return by throw vs. insertion of break). We thus classify both faults as (unconditional) control-flow target faults.
Our scheme comprises six control-flow and two data-flow fault classes. We use it to classify 488 faults from seven projects in the Defects4J dataset; the classification results are available at https://github.com/alexvdspuy/fcd4j. The most common individual class is one of the data-flow fault classes (definition fault); however, the control-flow classes are overall more common. Similar to repair classes [2], our fault classes are not disjoint since a single fault can have characteristics of multiple classes. The majority of the faults are thus assigned one to three of the eight classes.
II Flow Graph-Based Classification
II-A Flow graphs
We base our classification on flow graph representations of programs. Specifically, we consider the control-flow graph (CFG) and data-flow graph (DFG) as a single structure $G=(V,E_{C},E_{D})$, where the vertices $V$ correspond to the program statements, the CFG edges $E_{C}$ correspond to changes in the program counter, and the DFG edges $E_{D}$ represent associations between variable writes or defs and their references or uses. The flow graph construction, with explicit def-use association edges, is fairly standard [6, 7]. For the classification we assume that ($i$) jumps (i.e., break-, continue-, return-, and throw-statements) are represented by CFG edges only, and not by vertices, and ($ii$) method calls are represented by a vertex modeling the parameter passing and a CFG edge modeling the jump. We list our fault classes in §II-B to II-C.
II-B Control-flow faults
II-B1 Unconditional control-flow target faults
Such faults occur when at least one CFG edge has the wrong target and the source of the faulty edge has no other outgoing edge, i.e., is not a predicate node. Any run that reaches the source will thus traverse the faulty edge and inevitably enter an error state, but not necessarily fail. Faulty edges can correspond to wrong statement orders, or to missing, extraneous, or wrong jumps and method calls. Furthermore, the flow graph must already have the correct set of vertices, unless the fault occurs in combination with other faults.
Statement order faults
Such faults occur when a sequence of consecutive edges connect multiple vertices in the wrong order, but the set of source and target vertices of these edges is correct; a repair rearranges the edges’ sources and targets, which corresponds to reordering the statements. Closure-$13$ shows an example:
+ traverse(c);
Node next = c.getNext();
+ traverse(c);
c = next;
Listing 1: Statement order fault (Closure-$13$)
Jump faults
A jump fault occurs if the target of a single edge is wrong; this edge can connect two adjacent statements, representing the normal control flow, or span a larger gap, representing a jump statement. In both cases the repair effectively “re-wires” this edge. For example, Chart-$18$ contains a wrong jump, where the return-statement induces a CFG edge with the wrong target; replacing it with a throw-statement induces an edge with the correct target:
if (index < 0) {
- return;
+ throw new UnknownKeyException("The key (" + key
+ + ") is not recognised.");
Listing 2: Jump fault—wrong jump target (Chart-$18$)
In contrast, Lang-$62$ contains a missing jump fault where the control flow incorrectly connects the assignment to the next case-label, and the insertion of the break-statement repairs this. Similarly, removing a faulty extraneous jump restores the normal control flow.
case 'x' : {
entityValue = Integer.parseInt(entityContent
.substring(2), 16);
+ break;
Listing 3: Jump fault—missing jump (Lang-$62$)
Method call faults
The similarity between jumps and method calls also extends to faults. The only small difference is that the repair of a missing method call with arguments also requires inserting the vertex that models the parameter passing; this implies that such faults are also definition faults (§II-C1). Listings LABEL:listing:time22 and LABEL:listing:math64 show a wrong resp. missing method call (cf. Listings LABEL:listing:chart18 and LABEL:listing:lang62 for the corresponding jump faults).
- this(duration, null, null);
+ super();
Listing 4: Method call fault—wrong method (Time-$22$)
if (maxCosine <= orthoTolerance) {
// convergence has been reached
+ updateResidualsAndCost();
Listing 5: Method call fault—missing call (Math-$64$)
II-B2 Control-flow predicate faults
Control-flow predicates (i.e., the conditions of conditionals and loops) determine which CFG paths a program follows. A control-flow predicate fault occurs if at least one of these paths is incorrect. This can be caused by a fault in the condition itself, missing/extraneous predicate vertices guarding otherwise correct existing vertices, or entire predicate blocks that are missing.
In contrast to unconditional jump target faults, a program with a control-flow predicate fault may reach the fault location without entering an erroneous state (e.g., a run of the fault entry of Lang-$55$ shown in LABEL:listing:lang55, in which the inserted predicate would evaluate to true, does not cause an error).
Predicate vertex faults
This class captures faults where the CFG structure is correct, but the label of a predicate vertex (i.e., the condition) is wrong. For example, Chart-$1$ contains a predicate vertex fault, but if the faulty vertex is also the target of a faulty DFG edge (i.e., references a wrong variable), it is also classified as a data-flow fault.
- if (dataset != null) {
- if (dataset == null) {
return result;
Listing 6: Predicate vertex fault (Chart-$1$)
Predicate existence faults
This class captures faults where the CFG has an extraneous (resp. missing) predicate node guarding existing vertices, leading to paths that execute the guarded vertices in too few (resp. many) paths. For example, Lang-$55$ contains a missing if-guard fault:
+ if(this.runningState == STATE_RUNNING) {
stopTime = System.currentTimeMillis();
this.runningState = STATE_STOPPED;
Listing 7: Predicate existence fault—missing if (Lang-$55$)
Predicate block faults
In contrast to predicate existence faults, predicate block faults capture faults where the CFG is missing not only the predicate vertex but also the entire guarded block. Closure-$85$ contains such a fault:
Node next = ControlFlowAnalysis.computeFollowNode(n);
+ while (next != null && next.getType() == Token.BLOCK)
+ if (next.hasChildren())
+ next = next.getFirstChild();
+ else
+ next = computeFollowing(next);
return next;
Listing 8: Predicate block fault—missing while (Closure-$85$)
II-C Data-flow faults
Data-flow faults occur when the value written into a memory location does not match the requirements of a subsequent read from that location. This can be caused by faulty or missing definitions, or by faulty uses.
II-C1 Definition faults
Such faults occur at the source of a DFG edge. We distinguish between a wrong write value (i.e., the wrong value is written to the correct variable, see LABEL:listing:closure97), wrong write location (i.e., the correct value is written to the wrong variable), or a missing definition.
+ long lvalLong = lvalInt & 0xffffffffL;
- result = lvalInt $>$ rvalInt;
+ result = lvalLong $>$ rvalInt;
Listing 9: Definition fault—wrong write value (Closure-$97$)
II-C2 Use faults
Such faults occur at the target of a DFG edge, where the variable is read from the wrong location. Jsoup-57 contains a faulty use of attributes instead of it:
- attributes.remove(attrKey);
+ it.remove();
Listing 10: Use fault (Jsoup-$57$)
III Dataset Construction Methodology
We use Defects4J [8], a repository of real-world Java projects, as our data source. We classify the 488 faults from seven Defects4J projects (see Section III) with manual thematic analysis, similar to the process described by Sobreira et al. [2]. Each fault is inspected in terms of the source and repaired program, as well as failing test cases. The fault is then assigned the appropriate fault classes from our classification scheme. The first author of this paper performed manual classification, which was then reviewed by the second author, and another round of classification was done.
TABLE I: Project characteristics [7, 8, 9] and fault class frequency (maximum per project bolded)
{tblr}
[…]
Project KLOC order jump call pred guard block def use Avg per fault CF DF CF/DF All
[…]
Closure $90$ $6$ $9$ $37$ 4242{42}42 $53$ $30$ $\mathbf{81}$ $6$ $1.98\pm 1.03$ $52$ $14$ $67$ $133$
Lang $22$ $2$ $18$ $11$ $11$ $31$ $11$ $\mathbf{43}$ $8$ $2.08\pm 0.86$ $20$ $12$ $33$ $65$
Math $85$ $1$ $17$ $29$ $24$ $36$ $18$ $\mathbf{83}$ $20$ $2.15\pm 1.04$ $22$ $27$ $57$ $106$
Mockito $11$ 00 $3$ $15$ $2$ $16$ $6$ $\mathbf{28}$ $4$ $1.95\pm 0.86$ $10$ $4$ $24$ $38$
Time $28$ $3$ $9$ $6$ 77{7}7 $17$ $3$ $\mathbf{21}$ $5$ $2.63\pm 1.61$ $6$ $4$ $17$ $27$
[dotted] Jsoup $10$ $3$ $4$ $36$ $14$ $30$ $23$ $\mathbf{59}$ $6$ $1.88\pm 0.85$ $33$ $14$ $46$ $93$
[…]
We consider the difference in the flow graphs of the faulty and fixed code to manually assign fault classes according to the intrinsic fault nature. We infer the nature of the fault from the program context, taking into account the repair semantics. However, we investigate alternative repairs and compare all repair flow graphs to the faulty flow graph—the applied repair does not drive our classification.
We note that Sobreira et al. [2] were unable to classify 22 of the 395 faults in terms of their repair patterns; we are able to classify these faults in terms of our classification scheme. We also consider the Jsoup project, which has not been classified before and as such is unbiased by other classifications.
III-A Limitations
We consider, similarly to Sobreira et al. [2], each Defects4J fault entry as a single fault that may display characteristics of multiple classes. However, certain entries do contain multiple independent faults (e.g., Chart-$18$). For comparability and compatibility with the repair-based classification [2], we do not distinguish between multi-faults within an entry.
Several of the Defects4J patches combine the repair with refactoring (see LABEL:listing:math46). This biases a repair-based classification and may inhibit our fault classification.
IV Dataset Statistics
Section III shows the frequency of each fault class, in terms of the number of faults it is assigned to, over the 488 classified Defects4J faults. We see that definition fault (data-flow) is the most common fault class for each project. On the other hand, unconditional statement order fault (control-flow) is the least common class. Overall, the total control-flow classifications clearly outweigh the total data-flow classifications, despite individual control-flow classes being less frequently assigned.
We also see the distribution of the number of fault classes by the average number assigned per fault. All faults are assigned, on average, around two fault classes, with Time displaying the highest average ($2.63$). Furthermore, standard deviations of around $1.0$ show that faults are generally assigned between one and three fault classes. Figure 0(a) shows the co-assignments between our fault classes. Each cell $i,j$ shows the percentage of faults assigned fault class $i$ that are also assigned fault class $j$. There is high agreement between several classes, especially with definition fault. Due to the discussed distribution, faults are often assigned more than one class, and definition fault is the most frequent class.
Because the fault classes are not disjoint, we consider disjoint fault types, where a fault is assigned only control-flow classes (pure CF), only data-flow classes (pure DF), or both (CF/DF). The majority of faults display a mixture of control- and data-flow characteristics.
Sobreira et al. [2] investigate 395 Defects4J faults; however, they focus on the repair actions and patterns, rather than fault types. The repair actions are very fine-grained, since they distinguish between addition, removal, and modification in nearly all syntactic categories. This leads to 28 different classes. The nine repair patterns are more abstract, and thus closer to our classification. However, some repair pattern categories are not useful for our purposes. For example, single line does not give much information about the fault itself—the fault is fixed in one line, but this is not the only possible fix. Figure 0(b) shows the co-assignments of seven repair classes from [2], which show fewer agreements than our own. However, we argue that certain of their repair classifications are biased by the fix, and do not accurately represent the underlying fault. Consider LABEL:listing:lang7 and LABEL:listing:math46:
- if (str.startsWith("--")) {
- return null;
Listing 11: Conditional block repair—removal (Lang-$7$[2]
- return isZero ? NaN : INF;
+ return NaN;
Listing 12: Unwraps-from repair—if-else (Math-$46$[2]
The underlying fault in LABEL:listing:lang7 is not the conditional, but the return-statement; removing only the return would also be a valid fix. We instead classify this as an unconditional jump fault. The unwraps-from classification in LABEL:listing:math46 is biased by the fix, when the fault is a definition and use fault (referencing and returning INF instead of NaN). Figure 0(c) shows the percentage correlation between our assigned fault classes with the repair-based classes [2], and Figure 0(d) shows the reverse. We see several one-sided agreements between the classifications, with (cond block, guard) and (expr fix, pred) the only pairs to show two-sided majority agreement. This illustrates that the two classifications are distinct. Figures 0(a) and 0(d) show that both fault and repair class assignments often correlate with the definition fault class, implying that these faults and fixes often focus on definitions.
(a) Fault vs. fault
(b) Repair vs. repair
(c) Repair vs. fault (by fault)
(d) Repair vs. fault (by repair)
Figure 1: Agreement matrices between classes (our fault classes, described in §II; repair classes [2])
V Dataset Usage and Future Research Questions
Our fault classification scheme is useful for its exploration of intrinsic fault characteristics, instead of repairs. We suggest the use of our scheme and classified faults to investigate datasets similar to Defects4J and to reduce debugging costs.
We have already considered the research question “Are spectrum-based fault localization (SBFL) techniques more effective for certain fault types?” Our preliminary results show that SBFL performs worse on pure control-flow faults, likely because such faults often involve missing code, which is difficult to localize. SBFL techniques suitable to specific fault types may be developed, then tested on our classified faults and compared with existing techniques. We can also investigate the research question, “Can the classification of an unknown fault be predicted from SBFL performance?”
Other possible applications include automated program repair. For example, we can use the fault class frequency information to guide automated program repair to attempt fixes in order of fault class frequency; if we can classify individual faults on-the-fly, we could even tailor individual repair operations to the specific fault classes.
VI Related Datasets
Goodenough and Gerhart [10] give phase-based error classification (construction, specification, design, requirements), but this is too abstract for our purposes. They also classify program errors into missing control-flow paths, inappropriate path selection, and inappropriate or missing action. These classes are similar to some of our fault classes, but the classification does not account for all faults found in practice.
Duncan and Robson [11] focus on faults in C programs. Their main classes are problematic construct, multiple file usage, and portability issues. However, some of the faults are overly specific (e.g., confusion of & and &&) or not applicable to our dataset (e.g., scanf with wrong types).
Pan et al. [4] describe a classification scheme of 27 bug fix patterns. The scheme is quite fine-grained, e.g., change of if condition expression and change of loop predicate form different classes, despite both being predicate faults. However, since certain fault characteristics can be deduced from the fix patterns, the schema has found use and adoption. Nath et al. [3] use these classes as a fault classification scheme for Java, while Lucia et al. [5] limit their scheme to eight categories, six of which are taken from Pan et al. [4], with two categories added (incl. others). However, predicate faults remain divided between conditionals and loops.
DiGiuseppe and Jones [1] investigate fault localization performance wrt. four fault taxonomies, of which three are quite coarse, with only between four and five fault types. The single fine-grained taxonomy, taken from Hayes et al. [12], is a generic fault taxonomy and includes classes such as visual stimulation and documentation, which are typically not useful in bug datasets. DiGiuseppe and Jones find no correlation between their classes and localization effectiveness; however, our preliminary results show that such a correlation likely exists for our classification, indicating it is more suitable for such applications.
VII Conclusion
In this paper, we introduce a novel fault classification scheme based on program flow graphs. We classify 488 faults from seven real-world projects in the Defects4J repository. The distribution of the number of fault classes assigned per fault is generally similar between different projects; the average is around two classes assigned per fault, but faults are generally assigned between one and three classes. The fault class frequency over all faults shows that the definition fault (data-flow) is the most frequent fault class with 330 faults classified as such. The unconditional statement order fault (control-flow) is the least frequent, with only 16 faults classified as such. However, pure data-flow faults are the least common in the dataset ($16.6\%$), compared with pure control-flow ($31.6\%$) and mixed control- and data-flow ($51.8\%$) faults.